OpenStudy (anonymous):

let f(x)=lxl . find f(g(x)) and g(f(x)) for each given function . answer this please g(x)=x+3 & g(x)=2x

OpenStudy (zarkon):

if \[g(x)=x+3\] \[f(g(x))=f(x+3)=|x+3|\] \[g(f(x))=g(|x|)=|x|+3\]

OpenStudy (anonymous):

thank youuu !!

OpenStudy (zarkon):

can you do the other?

OpenStudy (anonymous):

im going to try

OpenStudy (anonymous):

G(x)=2x f(g(x))=l2xl g(f(x))=2lxl

OpenStudy (anonymous):

thankss